# Amorphous silicon structures generated using a moment tensor potential and the activation relaxation technique nouveau

### v1

@@ I Introduction
a-Si 7, 8 is idealized as a continuous random network (CRN) 9, characterized by a local atomic environment akin to crystalline silicon 10, 11. It maintains a local tetrahedral crystalline environment without long-range order. Real a-Si, however, is believed to contain defects associated with its disordered nature at a level that depends on its preparation12, 13, 14, 15, 6. These include local coordination defects, medium-range ring defects and two-level systems–the latter contribute to noise in sensitive gravitational wave detectors and cause decoherence in quantum computers 15, 6. Variations in tetrahedral geometry and bond lengths also significantly affect a-Si’s performance. Controlling structural anomalies during preparation poses challenges from both an experimental and a theoretical perspective. Indeed, the structure of a-Si is highly influenced by preparation methods 16; for instance, chemical vapor deposition 17 and ion implantation 18, 19, 13 can introduce voids and coordination defects. Thermal annealing 18, 20, 13, 21 often refines these imperfections, enhancing material quality.
@@
Extensive theoretical research has focused on developing realistic structural models for a-Si 10, 22, 23, 24, 25, 26, 27. Various methods for simulating materials, including a-Si, can be categorized into dynamic and non-dynamic approaches 7, 8. Dynamic methods, traditionally based on density functional theory (DFT) 28 or semiempirical force fields 29, primarily use melt-quench processes. However, these methods face challenges: experimental a-Si cannot be produced by melt-and-quench, DFT simulations are limited in system size and quenching rate, and semi-empirical force fields are often compromised by limited accuracy and transferability issues.
@@
Non-dynamic methods include the Reverse Monte Carlo (RMC) approach 30, 31, 32, 33, 34. In RMC, the computer-generated a-Si structure is randomly modified in order for its simulated structure factor and radial distribution function g(r) to match experimental data. RMC’s main drawback is that it can yield multiple configurations for the same g(r) — in other word the solution is degenerate. To overcome this, hybrid RMC methods incorporate force field interactions, ensuring generated configurations are both consistent with experimental data and physically realistic. Variants like the force-enhanced atomic refinement 35, 36 and experimentally constrained molecular relaxation 37 utilize semi-empirical and DFT interactions, respectively. Other key non-dynamic methods include the Wooten-Winer-Weaire (WWW) bond-switching algorithm 10, which optimizes atomic configurations by minimizing energy, and the Activation-Relaxation Technique nouveau (ARTn) 22, 38, 39, 40, which combines activation and relaxation steps to discover low-energy structures. ARTn can be paired with various force fields and ab initio methods, making it suitable for applications that include relaxation of disordered systems and diffusion processes 41, 15, 42, 43, 44, 45, 46, 47, 48, 49.
@@
One of the key components in both dynamical and non-dynamical methods is the interaction potential. The ongoing challenge of balancing accuracy and computational cost in these methods has led to the development of machine learning interatomic potentials (MLIPs) 50, 51, 52. MLIPs combine descriptors of the atomic environment, regression methods, and DFT data. Many MLIPs frameworks have been developed, including artificial neural networks 50, 53, kernel-based methods 54, 55, and linear regression approaches 56, 57, among others. For a detailed description of MLIPs, including their implementation and testing, the reader is referred to the existing literature 52. In the case of a-Si, structures have been modeled using artificial neural networks 58, 59, 60 and Gaussian Approximation Potentials (GAP) 54, 55. In initial studies, the GAP potential was applied to a-Si systems containing between 512 and 4,096 atoms 24. Later, the GAP model was extended to larger systems of up to 100,000 atoms to investigate the the atomistic mechanisms underlying various structural transitions in disordered silicon 61. Other MLIPs, such as the Moment Tensor Potential 57 (MTP), have also been applied to a-Si 62. In these studies, melt-quench MD simulations were employed to generate a-Si structures.
@@
A natural next step is to couple ARTn 22, 38, 39, 40 with machine learning potential to generate a-Si models. The objective of this paper is to generate a-Si models using ARTn-ML, which could enhance our understanding of the a-Si model and open new avenues for further research. The manuscript is organized into three main sections. The Methods section outlines the principles of the ARTn method and details our input models, the ARTn-MTP simulation process, and the analytical tools used. The Results and Discussion Section presents findings from our analysis of the configurations generated by ARTn-MTP. Finally, the key insights and discoveries of our study are summarized in the conclusions.
@@ II.1 Activation Relaxation Technique nouveau
ARTn 22, 38 is a computational method used in the study of activated processes, particularly in the context of diffusion and transitions between states in condensed matter systems. ARTn is designed to simulate rare events that signify substantial changes in the microscopic structure of atomic systems, based on the efficient exploration of potential energy landscapes. ARTn employs an event-driven approach in which the system evolves by transitioning between different states or configurations. The concept is based on the understanding that within the configurational space, there is a distinct set of points representing saddle points and adjacent minima in the potential energy surface of the system. As the name implies, ARTn can be summarized as a two-step process 22: activation and relaxation. In the activation step, starting from a local energy minimum, the configuration is pushed toward a nearby first-order saddle point. In the relaxation step, the system is driven over the saddle point and then relaxed to a new energy minimum. The new configuration is accepted based on the Metropolis acceptance probability, which is given by $P_{\text{accept}}=\min\left(1,\exp\left(-\frac{\Delta E}{kT}\right)\right)$, where T is the fictitious temperature, set to 0.25 eV for our simulation, k is the Boltzmann constant, and $\Delta E$ is the energy difference between the old and new configurations 22, 63, 23. The search is restricted to paths that include only saddle points. The ARTn method has demonstrated its value in revealing rare events that conventional simulation methods may struggle to capture due to the slow dynamics and relaxation inherent in these systems. Detailed descriptions of the method can be found elsewhere 22, 38, 39, 64, 65.
@@ II.2 A moment tensor potential for silicon
In practical terms, ARTn serves as a saddle-point search engine, employing an interaction potential as both a force and energy calculator. The simulation’s accuracy is directly influenced by the selection of the force field. Previous studies have relied on semi-empirical force field 42, 22, 66, 15, 67 and ab initio methods 68, 69, 44. Here we employ a machine learning-based forcefield. This potential was developed to jointly describe Si, O, and SiO2 systems 62, and is an improvement over a previously reported Si MTP 70. This potential is based on the MTP framework 57, 71. A detailed description of the development and implementation of this MTP can be found in our previous work 62, 70.
@@
216-FEAR Randomly filled 216 Force hence atomic refinement 36
@@
512-MD Randomly filled 512 MD simulation using MTP potential
512-RMC Randomly filled 512 Reverse monte carlo simulation 72, 24
512-INV Randomly filled 512 Invariant environment refinement technique 72, 24
512-FEAR Randomly filled 512 Force hence atomic refinement 36
512-WWW Randomly filled 512 Bond switching algorithm 72, 24
@@
100k-GAP18 Unknown 100000 MD simulation using GAP potential 61
1000-aSi-SW Amorphous configuration 22, 38, 68 1000 ARTn runs using MTP potential
@@
Fig. 1: Key structural elements characterizing a-Si: (a) bond angle ($\theta$), coordination defects such as dangling bonds (3-fold bonded atoms) and floating bonds (5-fold bonded atoms), and ring structures, including 4-, 5-, 6-, and 7-membered rings. (b) A 100% 4-fold fully coordinated amorphous structure comprising 216 atoms, generated by ARTn-MTP simulation , with 14.40 atoms in a crystalline environment. The atoms shown in yellow within the box are in an amorphous environment, while those colored differently are in a crystalline environment, specifically cubic or hexagonal diamond.
II.3 Protocol for input models preparation and simulation details
@@
We prepared models containing 216, 512, 1000, and 4096 atoms. For the model with 216 atoms, we used two different simulation boxes corresponding to densities of 2.20 g/cm³ (216-R1) and 2.28 g/cm³ (216-R2), respectively. These two boxes were randomly filled with an overlap distance (initial atomic separation) of 2.3 Å for 2.20 g/cm³ and 2 Å for 2.28 g/cm³. For the 512-atom model, we employed two simulation boxes randomly filled with atoms, using the same densities as those applied to the 216-atom model. The first 512-atom model had a density of 2.20 g/cm³ (512-R1) with an initial overlap of 2.3 Å. The second 512-atom model (512-R2) along with the 1000-atom model (1000-R) were randomly filled with a density of 2.28 g/cm³ and an initial overlap of 2 Å. The search for events during ARTn-MTP simulations was initiated using configurations that were already at local minima. Herein, the ARTn simulations begins by applying random displacements to the chosen central atom and its neighbors, based on a user-defined local cut-off radius. Once the displacements are introduced, the algorithm searches for convergence to a saddle point, guided by the lowest curvature. After locating the saddle point, the minimization process begins to find the new minimum. The minimization is carried out using either the Steepest Descent (SD) or the Fast Inertial Relaxation Engine (FIRE) algorithm. In all ARTn-MTP simulations, we used a fictitious temperature of 0.25 eV in the Metropolis accept-reject criterion. Each simulation started with event searches initiated with a local cut-off radius of 3 Å, which the algorithm used to identify the region around the selected central atom which will be deformed. This cut-off radius was progressively increased to a maximum cut-off 5.5 Å in order toexpand the local zone of initial deformaton when the ARTn moves no longer led to energy decrease. The coordination defects level was monitored during the simulations. For coordination defect levels above 5 %, the full list of atoms was used as potential central atoms for ARTn moves. For defect levels at or below 5 %, only atoms with a non-four-fold coordination were allowed to serve as the central atom in ARTn moves.
II.4 Structural analysis
@@
We employ a set of criteria to assess the quality of our amorphous structures. These criteria are derived from considerations of short-range structures, coordination defects, and intermediate-range order. Understanding these different structural aspects is crucial in the study of materials properties, as they influence the physical, chemical, and mechanical behavior of materials. We conduct coordination analysis using OVITO 73 with a bond-length cutoff of 2.85 Å. The analysis of the ring size distribution, providing insights into intermediate-range order, is performed using the R.I.N.G.S code 74 based on the King criteria. We employed the ISAACS code 75 for the calculation of the structure factor while the radial distribution function and angular distribution function were determined using LAMMPS 76 through pseudo simulations.
We also measure the amount of crystallinity in the models. Here, crystallity refers to environments that are cubic or hexagonal diamond to second neighbour, all atoms at the center of this environment or within the first- and second-neighbour shell of a perfectly atom are considered as having crystallinity. We use method developed by Ref. 77 as implemented in OVITO 73.

### v2

@@ I Introduction
a-Si 7, 8 is idealized as a continuous random network (CRN) 9, characterized by a local atomic environment akin to crystalline silicon 10, 11. It maintains a local tetrahedral crystalline environment without long-range order. Real a-Si, however, is believed to contain defects associated with its disordered nature at a level that depends on its preparation12, 13, 14, 15, 6. These include local coordination defects, medium-range ring defects and two-level systems–the latter contribute to noise in sensitive gravitational wave detectors and cause decoherence in quantum computers 15, 6. Variations in tetrahedral geometry and bond lengths also significantly affect a-Si’s performance. Controlling structural anomalies during preparation poses challenges from both an experimental and a theoretical perspective. Indeed, the structure of a-Si is highly influenced by preparation methods 16; for instance, chemical vapor deposition 17 and ion implantation 18, 19, 13 can introduce voids and coordination defects. Thermal annealing 18, 20, 13, 21 often refines these imperfections, enhancing material quality. Additionally, a recent study 22 has raised an important issue that questions the definition of the true model of a-Si. This study explored the footprints of paracrystallinity in amorphous silicon using molecular dynamics simulations with a machine learning potential 22, 23 similar to the one employed herein. As reported by the authors, while optimal models of a-Si are traditionally based on the continuous random network model, some groups have proposed that real a-Si could be better described as ’paracrystalline’. However, since both optimal types of structure conflict with experimental observations 24, 25, 26, 27, the authors have further investigated the paracrystalline model, concluding that realistic experimental models of a-Si may contain some degree of crystalline grain traces. The observed degree of crystallinity is likely influenced by the preparation method—i.e. MD. Notably, while the degree of crystallinity was examined, its relationship to coordination defects was not. In other words, the central question remains how to balance crystallinity and coordination defects when preparing a-Si. For instance, a high cooling rate in melt-quench simulations will likely lead to a higher number of coordination defects and lower crystallinity, while a low cooling rate will most likely yield fewer coordination defects and higher crystalline-like environments.
@@
Extensive theoretical research has focused on developing realistic structural models for a-Si 10, 28, 29, 30, 31, 32, 33. Various methods for simulating materials, including a-Si, can be categorized into dynamic and non-dynamic approaches 7, 8. Dynamic methods, traditionally based on density functional theory (DFT) 34 or semiempirical force fields 35, primarily use melt-quench processes. However, these methods face challenges: DFT simulations are limited in system size and quenching rate, and semi-empirical force fields are often compromised by limited accuracy and transferability issues.
@@
Non-dynamic methods include the Reverse Monte Carlo (RMC) approach 36, 37, 38, 39, 40. In RMC, the computer-generated a-Si structure is randomly modified in order for its simulated structure factor and radial distribution function g(r) to match experimental data. RMC’s main drawback is that it can yield multiple configurations for the same g(r) — in other word the solution is degenerate. To overcome this, hybrid RMC methods incorporate force field interactions, ensuring generated configurations are both consistent with experimental data and physically realistic. Variants like the force-enhanced atomic refinement 41, 42 and experimentally constrained molecular relaxation 43 utilize semi-empirical and DFT interactions, respectively. Other key non-dynamic methods include the Wooten-Winer-Weaire (WWW) bond-switching algorithm 10, which optimizes atomic configurations by minimizing energy, and the Activation-Relaxation Technique nouveau (ARTn) 28, 44, 45, 46, which combines activation and relaxation steps to discover low-energy structures. ARTn can be paired with various force fields and ab initio methods, making it suitable for applications that include relaxation of disordered systems and diffusion processes 47, 15, 48, 49, 50, 51, 52, 53, 54, 55.
@@
One of the key components in both dynamical and non-dynamical methods is the interaction potential. The ongoing challenge of balancing accuracy and computational cost in these methods has led to the development of machine learning interatomic potentials (MLIPs) 56, 57, 58. MLIPs combine descriptors of the atomic environment, regression methods, and DFT data. Many MLIPs frameworks have been developed, including artificial neural networks 56, 59, kernel-based methods 60, 61, and linear regression approaches 62, 63, among others. For a detailed description of MLIPs, including their implementation and testing, the reader is referred to the existing literature 58. In the case of a-Si, structures have been modeled using artificial neural networks 64, 65, 66 and Gaussian Approximation Potentials (GAP) 60, 61. In initial studies, the GAP potential was applied to a-Si systems containing between 512 and 4,096 atoms 30. Later, the GAP model was extended to larger systems of up to 100,000 atoms to investigate the the atomistic mechanisms underlying various structural transitions in disordered silicon 67. Other MLIPs, such as the Moment Tensor Potential 63 (MTP), have also been applied to a-Si 68. In these studies, melt-quench MD simulations were employed to generate a-Si structures.
@@
A natural next step is to couple ARTn 28, 44, 45, 46 with machine learning potential to generate a-Si models. The objective of this paper is to generate a-Si models using ARTn-ML, which could enhance our understanding of the a-Si model and open new avenues for further research. The manuscript is organized into three main sections. The Methods section outlines the principles of the ARTn method and details our input models, the ARTn-MTP simulation process, and the analytical tools used. The Results and Discussion Section presents findings from our analysis of the configurations generated by ARTn-MTP. Finally, the key insights and discoveries of our study are summarized in the conclusions.
@@ II.1 Activation Relaxation Technique nouveau
ARTn 28, 44 is a computational method used in the study of activated processes, particularly in the context of diffusion and transitions between states in condensed matter systems. ARTn is designed to simulate rare events that signify substantial changes in the microscopic structure of atomic systems, based on the efficient exploration of potential energy landscapes. ARTn employs an event-driven approach in which the system evolves by transitioning between different states or configurations. The concept is based on the understanding that within the configurational space, there is a distinct set of points representing saddle points and adjacent minima in the potential energy surface of the system. As the name implies, ARTn can be summarized as a two-step process 28: activation and relaxation. In the activation step, starting from a local energy minimum, the configuration is pushed toward a nearby first-order saddle point. In the relaxation step, the system is driven over the saddle point and then relaxed to a new energy minimum. The new configuration is accepted based on the Metropolis acceptance probability, which is given by $P_{\text{accept}}=\min\left(1,\exp\left(-\frac{\Delta E}{kT}\right)\right)$, where T is the fictitious temperature, set to 0.25 eV for our simulation, k is the Boltzmann constant, and $\Delta E$ is the energy difference between the old and new configurations 28, 69, 29. The search is restricted to paths that include only saddle points. The ARTn method has demonstrated its value in revealing rare events that conventional simulation methods may struggle to capture due to the slow dynamics and relaxation inherent in these systems. Detailed descriptions of the method can be found elsewhere 28, 44, 45, 70, 71.
@@ II.2 A moment tensor potential for silicon
In practical terms, ARTn serves as a saddle-point search engine, employing an interaction potential as both a force and energy calculator. The simulation’s accuracy is directly influenced by the selection of the force field. Previous studies have relied on semi-empirical force field 48, 28, 72, 15, 73 and ab initio methods 74, 75, 50. Here we employ a machine learning-based forcefield. This potential was developed to jointly describe Si, O, and SiO2 systems 68, and is an improvement over a previously reported Si MTP 76. This potential is based on the MTP framework 63, 77. The MTP potential was trained using data obtained from DFT calculations with the generalized gradient approximation (GGA-PBE) as the exchange-correlation functional 78. The MTP training set did not include amorphous silicon (a-Si) configurations. As explained in Ref. 68, the MTP was validated against ab initio MD by performing a rapid melt-and-quench. Fast cooling (–8 ps) was followed by equilibration (10 ps) using either ab initio MD or MTP-MD. Both ab initio MD and MTP-MD led to a-Si models with very similar structural properties while repeating the same protocol with semi-empirical potentials led to substantially different results. The simulations presented herein can be seen as further validation and testing of this MTP. A detailed description of the development and implementation of this MTP can be found in our previous work 68, 76.
II.3 Nomenclature of the models
To name our models, including those from the literature, we used several techniques. Some models were named based on their creation procedure, while others were named according to the simulation method and the year the model was generated. For our models, we used the following naming conventions. Five of our models were named 216-R1, 216-R2, 512-R1, 512-R2, and 1000-R. Here, R stands for randomly generated initial structure, meaning the simulation box was randomly filled with atoms. When there are multiple models with the same number of atoms, the first model is named R1, and the second is named R2. If there is only one model for a given number of atoms, it is simply named R, as in 1000-R. One of our models was initially randomly filled and then pre-relaxed with MD simulation, so it was named 4096-R-MD. The models named 216-MD1, 216-MD2, 512-MD1, 512-MD2, 1000-MD, and 4096-MD refer to configurations that were entirely simulated using Molecular Dynamics (MD) with the ML (MTP) potential. The initial configurations are the same as those used for 216-R1, 216-R2, 512-R1, 512-R2, 1000-R, and 4096-R-MD, meaning the initial setups were randomly generated as described earlier. We found these input configurations to be a good starting point, as opposed to starting from a crystalline state and melting it. Models from the literature retain their original preparation methods, as indicated in the last column of Table I. For example, 216-FEAR indicates that the a-Si model was prepared using Force-enhanced atomic refinement, and 100k-GAP18 refers to a model with 100,000 atoms of a-Si, prepared using the Gaussian Approximation Potential (GAP) developed in 2018 for silicon materials.
@@ II.3 Nomenclature of the models
216-MD1 Randomly filled 216 MD simulation using MTP potential
216-MD2 Randomly filled 216 MD simulation using MTP potential
216-FEAR Randomly filled 216 Force-enhanced atomic refinement 42
@@
512-MD1 Randomly filled 512 MD simulation using MTP potential
512-MD2 Randomly filled 512 MD simulation using MTP potential
512-RMC Randomly filled 512 Reverse monte carlo simulation 79, 30
512-INV Randomly filled 512 Invariant environment refinement technique 79, 30
512-FEAR Randomly filled 512 Force-enhanced atomic refinement 42
512-WWW Randomly filled 512 Bond switching algorithm 79, 30
@@
4096-MD Randomly filled + MD runs 4096 MD simulation using MTP potential
@@
100k-GAP18 Unknown 100000 MD simulation using GAP potential 67
1000-aSi-SW Amorphous configuration 28, 44, 74 1000 ARTn runs using MTP potential
@@
Fig. 1: Key structural elements characterizing a-Si: (a) bond angle ($\theta$), coordination defects such as dangling bonds (3-fold bonded atoms) and floating bonds (5-fold bonded atoms), and ring structures, including 4-, 5-, 6-, and 7-membered rings. (b) A 100% 4-fold fully coordinated amorphous structure comprising 216 atoms, generated by ARTn-MTP simulation , with 14.4 % of atoms in a crystalline environment. The atoms shown in yellow within the box are in an amorphous environment, while those colored differently are in a crystalline environment, specifically cubic or hexagonal diamond.
II.4 Protocol for input models preparation and simulation details
@@
We prepared models containing 216, 512, 1000, and 4096 atoms. For the model with 216 atoms, we used two different simulation boxes corresponding to densities of 2.20 g/cm³ (216-R1) and 2.28 g/cm³ (216-R2), respectively. These two boxes were randomly filled with an overlap distance (initial atomic separation) of 2.3 Å for 2.20 g/cm³ and 2 Å for 2.28 g/cm³. For the 512-atom model, we employed two simulation boxes randomly filled with atoms, using the same densities as those applied to the 216-atom model. The first 512-atom model had a density of 2.20 g/cm³ (512-R1) with an initial overlap of 2.3 Å. The second 512-atom model (512-R2) along with the 1000-atom model (1000-R) were randomly filled with a density of 2.28 g/cm³ and an initial overlap of 2 Å. The search for events during ARTn-MTP simulations was initiated using configurations that were already at local minima. Herein, the ARTn simulations begins by applying random displacements to the chosen central atom and its neighbors, based on a user-defined local cut-off radius. Once the displacements are introduced, the algorithm searches for convergence to a saddle point, guided by the lowest curvature. After locating the saddle point, the minimization process begins to find the new minimum. The minimization is carried out using either the Steepest Descent (SD) or the Fast Inertial Relaxation Engine (FIRE) algorithm. In all ARTn-MTP simulations, we used a fictitious temperature of 0.25 eV in the Metropolis accept-reject criterion.The fictitious temperature is not a thermodynamic temperature of our system as we use ARTn here was a way to minimize total energy through activated events. Since no thermodynamical ensemble is therefore defined, this parameter serves as an artificial tool to aid in sampling configurations and achieving better convergence in the simulation. Each simulation started with event searches initiated with a local cut-off radius of 3 Å, which the algorithm used to identify the region around the selected central atom which will be deformed. The cut-off radius defines a spherical region around the central atom, within which both the selected central atom and its neighboring atoms are displaced during each event. This cut-off radius was progressively increased to a maximum cut-off 5.5 Å in order to expand the local zone of initial deformaton when the ARTn moves no longer led to energy decrease. The coordination defects level was monitored during the simulations. For coordination defect levels above 5 %, the full list of atoms was used as potential central atoms for ARTn moves. For defect levels at or below 5 %, only atoms with a non-four-fold coordination were allowed to serve as the central atom in ARTn moves. For comparison purposes, we used the same initial (input) configurations as those for the ARTn-MTP simulations to run MD simulations with the MTP potential. For each MD model—216-MD1, 216-MD2, 512-MD1, 512-MD2, 1000-MD, and 4096-MD—the configuration was first heated to 2500 K over 30 ps, followed by equilibration at this temperature for 1 ns. The system was then cooled at a rate of 1 K/ps, and finally, the configuration was equilibrated at 300 K for 150 ps using the isothermal-isobaric ensemble (NPT).
II.5 Structural analysis
@@
We use a set of criteria to assess the quality of our amorphous structures, based on short-range order, coordination defects, intermediate-range order, and the presence of crystalline grains, as illustrated in Fig. 1 a). Understanding these different structural aspects is crucial in the study of materials properties, as they influence the physical, chemical, and mechanical behavior of materials. We conduct coordination analysis using OVITO 80 with a bond-length cutoff of 2.85 Å. The analysis of the ring size distribution, providing insights into intermediate-range order, is performed using the R.I.N.G.S code 81 based on the King criteria. We employed the ISAACS code 82 for the calculation of the structure factor while the radial distribution function and angular distribution function were determined using LAMMPS 83 through pseudo simulations.
We also quantify the crystallinity in the models, as indicated by the colored atoms (excluding the yellow atoms) in Fig. 1 b. Here, crystallity refers to environments that are cubic or hexagonal diamond to second neighbour, all atoms at the center of this environment or within the first- and second-neighbour shell of a perfectly atom are considered as having crystallinity. We use method developed by Ref. 84 as implemented in OVITO 80. To verify the crystalline content of our models, we used Common Neighbor Analysis (CNA) in OVITO, which offers four operation modes. Two are relevant for amorphous systems: the conventional CNA with a fixed cutoff and the Adaptive CNA with a variable cutoff. The conventional CNA is better suited for crystalline systems, while the Adaptive CNA is designed for multiphase systems, which is more relevant for amorphous systems with crystalline seeds. Both methods were applied to our models, but they are limited to lattice systems other than diamond and hexagonal diamond. Therefore, we then used the "Identify Diamond Structure" mode in OVITO, an extension of Adaptive CNA, which automatically determines the optimal cutoff radius for each particle. Finally, we varied the cutoff in conventional CNA but observed no changes in crystallinity.